# Towards Understanding Why FixMatch Generalizes Better Than Supervised Learning

### v1

@@ Appendix G Proof for FlexMatch, FreeMatch, Dash, and SoftMatch
FlexMatch (Zhang et al., 2021a) differentiates itself from FixMatch by modifying the constant threshold $\tau$ to include an adaptive class-specific threshold $\beta_{t}(b)$ for each class $b$. Under our multi-view data assumption as defined in Definition 7, the data distribution for each class is the same. Consequently, as suggested by Claim 24 and Claim D.10 in Allen-Zhu & Li (2023), all classes progress at a similar rate during training. This uniformity over all classes allows us to standardize $\beta_{t}(b)=1$, $\forall b\in[k]$, thereby aligning the proof for FlexMatch with that of FixMatch.
For FreeMatch (Wang et al., 2022b) and SoftMatch (Chen et al., 2023), instead of applying a large constant threshold $\tau$ during the training process, they use an adaptive $\tau_{t}$ to involve more unlabeled data with correctly-predicted pseudo-label in the training of the network. Under our multi-view data assumption 7, the majority of the data in training dataset is of multi-view (with probability $1-\frac{1}{\mathrm{poly}(k)}$), so we only consider the network’s prediction confidence for multi-view data to determine $\tau_{t}$. We set the adaptive threshold $\tau_{t}$ as follows:
@@
where $\beta$ is the momentum parameter, $b=\operatorname*{arg\,max}_{i}\mathbf{logit}_{i}(F^{(t)},X)$, and $T_{0}=\Theta(\frac{k}{\eta\sigma_{0}^{q-2}})$. Here we do not consider the unsupervised loss term Eq. (5) before $T_{0}$-th iteration in our analysis, since the model is bad at generating correct pseudo-label at the initial phase of training. We use $\max$ function here to ensure the high quality of unlabeled data involved at each training step. After $T_{0}$-th iteration, according to Claim D.11 and Lemma D.22 in Allen-Zhu & Li (2023), the feature correlations increase to $\Lambda_{i}^{(t)}=\tilde{\Theta}(1)$ for $t\geq T_{0}$, while the off-diagonal correlations $\langle w_{i,r},v_{j,l}\rangle$ ($i\neq j$) keep small at the scale of $\tilde{O}(\sigma_{0})$. Denote $\Phi^{(t)}=\max_{i\in[k],l\in[2]}\Phi_{i,l}^{(t)}$, recall from Claim D.9 in Allen-Zhu & Li (2023), for every $X\in\operatorname*{arg\,max}_{X\in\mathcal{Z}_{u,m}}[\mathbf{logit}_{b}(F^{(t)},X)]$ with ground truth label $y$, we have $F_{j}^{(t)}(X)\leq 0.8001\Phi^{(t)}$ for $j\neq y$ and $F_{y}^{(t)}(X)\geq 0.9999\Phi^{(t)}$ with probability at least $1-e^{-\Omega(\log^{2}k)}$. Accordingly, we have $F_{y}^{(t)}(X)\geq\max_{j\neq y}F_{j}^{(t)}(X)+\tilde{\Theta}(1)$, which means that $F^{(t)}$ can correctly classify the unlabeled data with high probability (i.e., $b=y$). Therefore, when $t\geq T_{0}$, both the supervised loss Eq. (4) and unsupervised loss Eq. (5) take effect. Same as in Sec. 4, we use $\Phi_{i,l}^{(t)}$ here to monitor the feature learning process. For $(i,l)\in\mathcal{M}_{F}$, feature $v_{i,l}$ is partially learned during the first $T_{0}$ iterations in that $\Phi_{i,l}^{(T_{0})}=\tilde{\Theta}(1)<\Omega(\log(k))$, while feature $v_{i,3-l}$ is missed in that $\Phi_{i,3-l}^{(T_{0})}=\tilde{O}(\sigma_{0})\ll\tilde{\Theta}(1)$. Start from $T_{0}$, feature $v_{i,l}$ is continued to be better learned with the help of supervised loss and unsupervised loss until $\Phi_{i,l}^{(t)}\geq\Omega(\log k)$, and feature $v_{i,3-l}$ start to be learned with the help of unsupervised loss. We can analyze this feature learning process using a similar approach as in Sec. E. The key intuition for the extension of the proof of FixMatch to FreeMatch and SoftMatch is that by setting an adaptive confidence threshold, the learning process of unlearned features begin at $T_{0}$ instead of $T=O(\mathrm{poly}(k)/\eta)>T_{0}$ in FixMatch.
@@ Appendix H Effect of Strong Augmentation on Supervised Learning
In this section, we show why using strong augmentation with probabilistic feature removal effect, such as CutOut, in supervised learning (SL) has minimal alternation to the feature learning process. In SL, strong augmentation $\mathcal{A}(\cdot)$ is utilized at the start of training, before any feature has been effectively learned, corresponding to Phase I of SSL. According to Assumption 9, $\mathcal{A}(\cdot)$ randomly removes its semantic patches and noisy patches with probabilities of $\pi_{2}$ and $1-\pi_{2}$, respectively. Then for a single-view image, its only semantic feature is masked with probability $\pi_{2}$. For a multi-view image, one of the two features, $v_{i,1}$ or $v_{i,2}$ is masked with probabilities $\pi_{1}\pi_{2}$ and $(1-\pi_{1})\pi_{2}$, respectively. Thus, the size of single-view data in training dataset $\mathcal{Z}_{l}$ is increased, as $\mathcal{A}(\cdot)$ transfers $\pi_{2}N_{l,m}$ multi-view samples to single-view.
@@ Appendix I Comparison with Pioneering Work
Data augmentation operations like CutOut can help supervised learning (SL), but cannot improve as much as in deep SSL with limited labeled data. On STL-10 dataset with 40 labeled data, when we remove CutOut from SL, the test accuracy (%) does not drop a lot as shown in Table 8. In contrast, removing CutOut from the strong augmentation $\mathcal{A}(\cdot)$ in FixMatch’s unsupervised loss $L_{u}^{(t)}$ leads to a severe performance degradation.
@@
All the experiments are conducted on four RTX 3090 GPU (24G memory). Due to limited resources, we did not train the models for 1024 epochs with 1024 iterations per epoch (in total $2^{20}$ iterations) as in Sohn et al. (2020), but run 150 epochs with 2048 iterations per epoch (in total $307200$ iterations). According to our experimental results in Sec. 5, the test accuracy results of our training setting approximates the results obtained by Sohn et al. (2020). For the experiments on FixMatch, our code is based on Kim (2020); for all other experiments, our code is based on Wang et al. (2022a). Each individual SSL experiment requires 48 to 120 hours to complete on a single RTX 3090 GPU, depending on the model and dataset.
@@
Table 10: Complete hyperparameter setting.

### v2

@@ Appendix G Proof for FlexMatch, FreeMatch, Dash, and SoftMatch
FlexMatch (Zhang et al., 2021a) differentiates itself from FixMatch by modifying the constant threshold $\tau$ to include an adaptive class-specific threshold $\beta_{t}(b)$ for each class $b$. Under our multi-view data assumption as defined in Def. 7, the data distribution for each class is the same. Consequently, as suggested by Claim 24 and Claim D.10 in Allen-Zhu & Li (2023), all classes progress at a similar rate during training. This uniformity over all classes allows us to standardize $\beta_{t}(b)=1$, $\forall b\in[k]$, thereby aligning the proof for FlexMatch with that of FixMatch.
For FreeMatch (Wang et al., 2022b) and SoftMatch (Chen et al., 2023), instead of applying a large constant threshold $\tau$ during the training process, they use an adaptive $\tau_{t}$ to involve more unlabeled data with correctly-predicted pseudo-label in the training of the network. Under our multi-view data assumption Def. 7, the majority of the data in training dataset is of multi-view (with probability $1-\frac{1}{\mathrm{poly}(k)}$), so we only consider the network’s prediction confidence for multi-view data to determine $\tau_{t}$. We set the adaptive threshold $\tau_{t}$ as follows:
@@
where $\beta$ is the momentum parameter, $b=\operatorname*{arg\,max}_{i}\mathbf{logit}_{i}(F^{(t)},X)$, and $T_{0}=\Theta(\frac{k}{\eta\sigma_{0}^{q-2}})$. Here we do not consider the unsupervised loss term Eq. (5) before $T_{0}$-th iteration in our analysis, since the model is bad at generating correct pseudo-label at the initial phase of training. We use $\max$ function here to ensure the high quality of pseudo-label for unlabeled data involved at each training step. After $T_{0}$-th iteration, according to Claim D.11 and Lemma D.22 in Allen-Zhu & Li (2023), the feature correlations increase to $\Lambda_{i}^{(t)}=\tilde{\Theta}(1)$ for $t\geq T_{0}$, while the off-diagonal correlations $\langle w_{i,r},v_{j,l}\rangle$ ($i\neq j$) keep small at the scale of $\tilde{O}(\sigma_{0})$. Denote $\Phi^{(t)}=\max_{i\in[k],l\in[2]}\Phi_{i,l}^{(t)}$, recall from Claim D.9 in Allen-Zhu & Li (2023), for every $X\in\operatorname*{arg\,max}_{X\in\mathcal{Z}_{u,m}}[\mathbf{logit}_{b}(F^{(t)},X)]$ with ground truth label $y$, we have $F_{j}^{(t)}(X)\leq 0.8001\Phi^{(t)}$ for $j\neq y$ and $F_{y}^{(t)}(X)\geq 0.9999\Phi^{(t)}$ with probability at least $1-e^{-\Omega(\log^{2}k)}$. Accordingly, we have $F_{y}^{(t)}(X)\geq\max_{j\neq y}F_{j}^{(t)}(X)+\tilde{\Theta}(1)$, which means that $F^{(t)}$ can correctly classify the unlabeled data with high probability (i.e., $b=y$). Therefore, when $t\geq T_{0}$, both the supervised loss Eq. (4) and unsupervised loss Eq. (5) take effect. Same as in Sec. 4, we use $\Phi_{i,l}^{(t)}$ here to monitor the feature learning process. For $(i,l)\in\mathcal{M}$, feature $v_{i,l}$ is partially learned during the first $T_{0}$ iterations in that $\Phi_{i,l}^{(T_{0})}=\tilde{\Theta}(1)<\Omega(\log(k))$, while feature $v_{i,3-l}$ is missed in that $\Phi_{i,3-l}^{(T_{0})}=\tilde{O}(\sigma_{0})\ll\tilde{\Theta}(1)$. Start from $T_{0}$, feature $v_{i,l}$ is continued to be better learned with the help of supervised loss and unsupervised loss until $\Phi_{i,l}^{(t)}\geq\Omega(\log k)$, and feature $v_{i,3-l}$ start to be learned with the help of unsupervised loss. We can analyze this feature learning process using a similar approach as in Sec. E. The key intuition for the extension of the proof of FixMatch to FreeMatch and SoftMatch is that by setting an adaptive confidence threshold, the learning process of unlearned features begin at $T_{0}$ instead of $T_{1}=\mathrm{poly}(k)/\eta\gg T_{0}$ in FixMatch.
@@ Appendix H Effect of Strong Augmentation on Supervised Learning
In this section, we show why using strong augmentation with probabilistic feature removal effect, such as CutOut, in supervised learning (SL) has minimal alternation to the feature learning process. In SL, strong augmentation $\mathcal{A}(\cdot)$ is utilized at the start of training, before any feature has been effectively learned, corresponding to Phase I of SSL. According to Assumption 9, $\mathcal{A}(\cdot)$ randomly removes its semantic patches and noisy patches with probabilities of $\pi_{2}$ and $1-\pi_{2}$, respectively. Then for a single-view image, its only semantic feature is removed with probability $\pi_{2}$. For a multi-view image, one of the two features, $v_{i,1}$ or $v_{i,2}$ is removed with probabilities $\pi_{1}\pi_{2}$ and $(1-\pi_{1})\pi_{2}$, respectively. Thus, the size of single-view data in training dataset $\mathcal{Z}_{l}$ is increased, as $\mathcal{A}(\cdot)$ transfers $\pi_{2}N_{l,m}$ multi-view samples to single-view.
@@ Appendix I Comparison with Pioneering Work
Data augmentation operations like CutOut can help SL, but cannot improve as much as in deep SSL with limited labeled data. On STL-10 dataset with 40 labeled data, when we remove CutOut from SL, the test accuracy (%) does not drop a lot as shown in Table 8. In contrast, removing CutOut from the strong augmentation $\mathcal{A}(\cdot)$ in FixMatch’s unsupervised loss $L_{u}^{(t)}$ leads to a severe performance degradation.
@@
All experiments are conducted on four RTX 3090 GPUs (24GB memory). Due to resource limitations, we did not follow Sohn et al. (2020) by training for 1024 epochs with 1024 iterations per epoch ($2^{20}$ iterations in total). Instead, we run 150 epochs with 2048 iterations per epoch (307,200 iterations) for the standard setting in Sec. 5.1, and 50 epochs with 2048 iterations per epoch (102,400 iterations) for the same training dataset setting in Sec. 5.4. As shown in Sec. 5, our test accuracy results closely match those of Sohn et al. (2020). For FixMatch experiments, we base our implementation on Kim (2020), while all other experiments follow Wang et al. (2022a). Each SSL experiment takes 48 to 120 hours on a single RTX 3090 GPU, depending on the model and dataset.
@@
Table 10: Complete hyper-parameter setting.
@@
Appendix L Explanation of Multi-view Data Assumption
In this section, we make more detailed explanations of our multi-view data assumption Def. 1 by breaking it down and explain each part with specific examples from car images in ImageNet.
From Figure 1 we know that wheel and front light can be viewed as two discriminative semantic features of car, and each can be used independently for the learning model to make correct class prediction. In Figure 5, we give some examples of single-view data and multi-view data in car images that contains either only one of the two features or both. Then, we use them as examples to explain the multi-view data assumption Def. 1 in detail.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: The first single-view image contains only the front light feature, while the middle two multi-view images contain both wheel and front light features, and the last single-view image contains only the wheel feature.
In Def. 1, the data distribution $\mathcal{D}$ is composed of samples from the multi-view data distribution $\mathcal{D}_{m}$ with probability $1-\mu$, and from the single-view data distribution $\mathcal{D}_{s}$ with probability $\mu\!=\!\frac{1}{\mathrm{poly}(k)}$. In the context of car images, this implies that the majority of images are multi-view, containing both wheel and front light features, while a small fraction of images are single-view, containing only one of these features.
Then, Def. 1 defines $(X,y)\!\sim\!\mathcal{D}$ by first randomly selecting a label $y\!\in\![k]$ uniformly, and generate the data $X$ as follows:
(a) A set of noisy features $\mathcal{V}^{\prime}$ is sampled uniformly at random from $\{v_{i,1},v_{i,2}\}_{i\neq y}$, each with probability $s/k$. The complete feature set of $X$ is then defined as $\mathcal{V}(X)=\mathcal{V}^{\prime}\cup\{v_{y,1},v_{y,2}\}$, which includes both the noisy features $\mathcal{V}^{\prime}$ and the main features $\{v_{y,1},v_{y,2}\}$.
In the context of car images, (a) corresponds to the semantic features specific to cars (wheel and front light) being present in car images, along with noisy features from other classes, such as houses or trees in the background.
(b) For each $v\in\mathcal{V}(X)$, pick $C_{p}$ disjoint patches in $[P]$ and denote them as $\mathcal{P}_{v}(X)$. For a patch $p\in\mathcal{P}_{v}(X)$, we set $x_{p}=z_{p}v+``noises"\in\mathbb{R}^{d},$ where the coefficients $z_{p}\geq 0$ satisfy:
(b1) For "multi-view" data $(X,y)\in\mathcal{D}_{m}$, $\sum_{p\in\mathcal{P}_{v}(X)}z_{p}\in[1,O(1)]$ when $v\in\left\{v_{y,1},v_{y,2}\right\}$ and $\sum_{p\in\mathcal{P}_{v}(X)}z_{p}\in[\Omega(1),0.4]$ when $v\in\mathcal{V}(X)\setminus\left\{v_{y,1},v_{y,2}\right\}$.
(b2) For "single-view" data $(X,y)\in\mathcal{D}_{s}$, pick a value $\hat{l}\in[2]$ randomly uniformly as the index of the main feature. Then $\sum_{p\in\mathcal{P}_{v}(X)}z_{p}\in[1,O(1)]$ when $v=v_{y,\hat{l}}$, $\sum_{p\in\mathcal{P}_{v}(X)}z_{p}\in[\rho,O(\rho)]$ ($\rho=k^{-0.01}$) when $v=v_{y,3-\hat{l}}$, and $\sum_{p\in\mathcal{P}_{v}(X)}z_{p}=\frac{1}{\mathrm{polylog}(k)}$ when $v\in\mathcal{V}(X)\setminus\left\{v_{y,1},v_{y,2}\right\}$.
In the context of car images, (b1) indicates that multi-view data features a significant presence of two prominent class-specific semantic features, while the proportion of noisy features in the image is relatively small. For example, in the middle two multi-view car images shown in Figure 5, the wheel and front light features are more prominent than the background elements, such as houses and trees. (b2) indicates that single-view data contains only one prominent class-specific semantic feature, with the other semantic feature and noisy features being minimal. As shown in the first and last single-view car images in Figure 5, these images prominently display either the wheel or the front light, while the other semantic feature and noisy background elements are scarcely present.
In the context of car images, purely noisy patches correspond to regions such as road or sky patches that do not contain semantic information relevant to classification.
For a neural network capable of learning comprehensive semantic features after training–such as both the wheel feature and the front light feature for car images–can accurately predict both multi-view samples, such as the middle two images in Figure 5, and the single-view samples, such as the first and last images in Figure 5. However, if the network only learns partial semantic features, either the wheel or the front light feature, it will misclassify the single-view images that do not contain this feature, either the first or the last image in Figure 5 and result in inferior generalization performance.